OpenStudy (anonymous):

WILL GIVE MEDAL!!!!!!!!!!!! Estimate the quotient of seven and five over six divided by one and one over five using compatible numbers. What is the quotient? 4 9 8 2

OpenStudy (anonymous):

@Hero please help me, i don't understand

OpenStudy (anonymous):

anyone help

OpenStudy (anonymous):

please hero, i do not understand

hero (hero):

\[7 \frac{5}{6} =7 + \frac{5}{6}\]

OpenStudy (anonymous):

so i write it as an improper fraction?

hero (hero):

What does \(\dfrac{6}{6}\) equal?

OpenStudy (anonymous):

1 whole

hero (hero):

And 5/6 is close to 1 correct?

OpenStudy (anonymous):

yes

hero (hero):

So we can round 5/6 to 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aha

hero (hero):

Which means by way of rounding: \(7 + \dfrac{5}{6} \approx = 7 + 1 = 8\)

OpenStudy (anonymous):

yep

hero (hero):

Now, \(1 + \dfrac{1}{5}\)

hero (hero):

We will round that to the closest whole number

OpenStudy (anonymous):

okay

hero (hero):

To see this better, let's create a number line

OpenStudy (anonymous):

allright

OpenStudy (anonymous):

how old are you?

hero (hero):